how can $54000 be invested, part at 12%annual simple interest and the remainder at 6%.so that the interest earned by the two accounts will be eqaul
Answer by mananth(16949)

let investment @ 12% be x
and 6 % be y
x+y = 54000...................1
12%x=6*%y
multiply by 100
12x=6y
x= y/2
substitute x in (1)
y/2 + y =54000
multiply by 2
y+2y= 108000
3y=108000
/3
y=36000
$36,000 @ 6%
$18000 @ 12 %